ganeshie8 (ganeshie8):

A regular \(n-gon\) is inscribed in an unit circle. Let \(c_{ij}\) be the length of chord joining vertices \(i\) and \(j\). Show that \[\sum\limits_{i\ne j} {c_{ij}}^2 = n^2\] and \[\sum\limits_{i\ne j}\dfrac{1}{{c_{ij}}^2} = \dfrac{n^2-1}{12}\]

OpenStudy (freckles):

 how to show: \[c_{ij}=c_{jk}=c_{ki}=\sqrt{3}\]

OpenStudy (thomas5267):

\[ \operatorname{crd}(x)=2\sin\left(\frac{x}{2}\right)\\ \]

OpenStudy (thomas5267):

What would be the sum of n=3? \[ \begin{align*} \sum\limits_{i\ne j} {c_{ij}}^2 &= c_{12}+c_{13}+c_{23}?\\ &=c_{12}+c_{13}+c_{23}+c_{21}+c_{31}+c_{32}? \end{align*} \]

ganeshie8 (ganeshie8):

sry i think order doesn't matter freckles has correct interpretation it seems.. there are only \(\binom{n}{2}\) chords, not \(n^2\)

ganeshie8 (ganeshie8):

What would be the sum of n=3? \[ \begin{align*} \sum\limits_{i\ne j} {c_{ij}}^2 &= c_{12}+c_{13}+c_{23}?\\ \end{align*} \] i meant this, i see the ambiguity with my notation but i just don't know how to fix it...

OpenStudy (freckles):

 well this my interpretation for a 4-gon...\[c^2_{12}+c^2_{23}+c^2_{34}+c^2_{41}=4^2 \\ \text{ but this a a reg n-gon } \\ \text{ so } c_{12}=c_{23}=c_{34}=c_{41} \\ \text{ so this really becomes } 4 c_{12}^2=4^2 \\ \text{ and so we really just want to show } c_{12}^2=4 \text{ or that } c_{12}=2\]